Solve Algebraically using only one variable. Find three consecutive even integers such that the product of the first integer and the third integer is equal to four less than ten times the second integer.
Answer by stanbon(75887)

Find three consecutive even integers such that the product of the first integer and the third integer is equal to four less than ten times the second integer.
---------
Comment: Even numbers are always multiples of 2
----------
1st number: 2x
2nd number: 2x+2
3rd number: 2x+4
EQUATION:
(2x)(2x+4)=10(2x+2)-4
Divide thru by 4 to get:
x(x+2) = 5(x+1)-1
x^2+2x = 5x+4
x^2-3x-4=0
(x-4)(x+1)=0
Positive answer:
x=4
--------
1st number : 2x = 8
2nd number : 10
3rd number : 12
=========================
Cheers,
Stan H.